OpenStudy (anonymous):

Need help please!! A submicroscopic particle suspended in a solution has a volume 1.3um^3. What is this volume in liters? um = micrometers 1um = (10^-6) m

OpenStudy (freckles):

\[ 1 um =10^{-6} m\] do you know how many liters are equal to how many meters?

OpenStudy (freckles):

i'm sorry how many CUBIC meters equals how many liters?

OpenStudy (freckles):

\[1 m^3=10^3 L \text{ right ? }\]

OpenStudy (freckles):

\[1 um = 10^{-6 } m \\ \text{ cube both sides } \\ (1 um )^3=(10^{-6} m )^3 \\ 1 um^3=10^{-18} m^3 \]

OpenStudy (freckles):

do you know why I did this?

OpenStudy (freckles):

because we are working with cubic units

OpenStudy (freckles):

\[1.3 um^3 \cdot \frac{10^{-18} m^3}{1 um^3} \text{ this will cancel the } um^3 \text{ units } \] now we need to get rid of the m^3 units to do that we will also need m^3 on bottom and we know 1 m^3=10^3 L \[\text{so we know } \frac{ 10^3 L}{1 m^3 }=1\]

OpenStudy (freckles):

so we will also need to multiply that in

OpenStudy (freckles):

\[1.3 um^3 \cdot \frac{10^{-18} m^3}{1 um^3} \cdot \frac{10 ^3 L}{1 m^3}\]

OpenStudy (freckles):

so all the units will cancel there except the L units on top

OpenStudy (freckles):

which is what we wanted

OpenStudy (freckles):

so now just multiply the numbers on top and and you would divide by the numbers on bottom (but that is just one in this case :))

OpenStudy (anonymous):

ok thank you soooo much :)